OpenStudy (anonymous):

Express 27 as a product of it's prime factors using index notation.

OpenStudy (paxpolaris):

\[27 = 3\times 3 \times 3 =3^3\]

OpenStudy (skullpatrol):

divide by each prime in order as many times as possible

OpenStudy (skullpatrol):